the first side of a triangle is 5 more than 3 times the third side. the second side is 5 less than twice the third side. if the perimeter is 60cm, how long is each side?
Answer by bam878s(77)

Let x=length of third side, y=length of first side, and z=length of second side.
From the problem, we know
1. y=lenght of the first side= 5 more than 3 times the third side = 3x+5
2. z=length of the second side= 5 less than 2 times the third side = 2x-5
3. x=length of the third side.
4. The sum of these lengths, by definition is the perimeter.
We have,
perimeter=60= x + (3x+5) + (2x-5).
solve for x.
60=x+3x+5+2x-5
60=6x+5-5
60=6x
10=x
Now plug x back into your other equations to solve for the other sides.
y=length of first side=3x+5 = 3(10)+5 = 30+5 = 35
z=length of second side = 2x-5 = 2(10)-5= 20-5 = 15.
so
x=10
y=35
z=15.